OpenStudy (anonymous):

Showing all of your work, solve for x: 3(x + 2) = -5 -2(x -3)

OpenStudy (haleyelizabeth2017):

First, you need to distribute the 3 to the x and 2, and the -2 to the x and -3.

OpenStudy (anonymous):

ok give me a sec

OpenStudy (anonymous):

okay so it would turn into 3x+6

OpenStudy (haleyelizabeth2017):

That would be the left, correct. Now how about the right side? Ignore the -5 until after you are done

OpenStudy (anonymous):

-2x-6

OpenStudy (haleyelizabeth2017):

Close. Look at the negative signs.....a negative times a negative always equals a positive.

OpenStudy (anonymous):

yea so it would be -2x - 6

OpenStudy (haleyelizabeth2017):

-2*-3 is -6? I don't think so :P

OpenStudy (anonymous):

ohhhhhhh 6

OpenStudy (haleyelizabeth2017):

So now you have what?

OpenStudy (anonymous):

-2x 6

OpenStudy (haleyelizabeth2017):

I meant overall lol like the other side too :P

OpenStudy (anonymous):

3x+6=-5-2x6

OpenStudy (haleyelizabeth2017):

3x+6=-5-2x+6 you forgot the addition sign ;)

OpenStudy (haleyelizabeth2017):

So now you need to combine like terms on the right. What is -5+6?

OpenStudy (anonymous):

1?

OpenStudy (haleyelizabeth2017):

Correct! Now you just have -2x+1 on the right.

OpenStudy (haleyelizabeth2017):

So now, you want to get the variable on the right to the left side. Do you know what to do for that?

OpenStudy (here_to_help15):

Great explanation @haleyelizabeth2017 keep up the good work :)